OpenStudy (anonymous):

simplify in form of a +bi (2+3i)^-2 simplify in form of a +bi (2+3i)^-2 @Mathematics

OpenStudy (curry):

problem is typed right

OpenStudy (curry):

not typed*

OpenStudy (anonymous):

\[(2+3i)^-2\]

OpenStudy (anonymous):

that's root -2

OpenStudy (anonymous):

my guess is:\[\sqrt{2}+\sqrt{3i}, \sqrt{2}-i \sqrt{3}\]

OpenStudy (curry):

nope when u have negative it becomes a fraction

OpenStudy (nocipher):

(2+3i)^-2 = 1/(2+3i)^2 = 1/(-5 + 6i)

OpenStudy (curry):

ur square root answer would be right if the exponent was 1/2

OpenStudy (curry):

nocipher is wrong too

OpenStudy (anonymous):

ah I gotcha, Nocipher I want to understand how you got what you did. where did the -5+6 come from?

OpenStudy (curry):

its 1/(-5+12i)

OpenStudy (anonymous):

explain?

OpenStudy (nocipher):

Ah, Curry is right. Did the math in my head wrong

OpenStudy (curry):

u treat the problem like a binomial not like an expression without variables

OpenStudy (curry):

therefore use box method, except the asner will be denominator

OpenStudy (anonymous):

Man you guys make me feel soo dumb when it comes to math.

OpenStudy (nocipher):

Exponentiation of complex numbers by real numbers works just like it does for real numbers. The only case when exponentiation might get complicated is when you raise a number to a complex power.

OpenStudy (anonymous):

so \[=1/2^2 + 3i^2= 1/(2+3i)(2+3i)\]

OpenStudy (nocipher):

You can't distribute powers like that. Be careful (1/2^2) + 3i^2 = 1/4 - 3, not 1/((2+3i)(2+3i))

OpenStudy (curry):

4+12i+9i^2

OpenStudy (curry):

i^2 = 1

OpenStudy (curry):

so 4-9=-5

OpenStudy (curry):

-5+12i

OpenStudy (nocipher):

i^2 = -1, not 1

OpenStudy (anonymous):

yeah what I got was 1/12+12i, I'm supposed to clear the denominator too so I'm still workin.

OpenStudy (anonymous):

1/(12+12i)

OpenStudy (nocipher):

Ah, but you needed it in the form a+bi... so the answer we got was 1/(-5+12i). You can use the complex conjugate of the denominator, -5-12i, to further reduce your answer. So multiply 1/(-5+12i) by (-5-12i)/(-5-12i) and get (-5-12i)/169. Then in a+bi form that is (-5/169)+(-12/169)i.

OpenStudy (anonymous):

AHHH it is negative 5!

OpenStudy (anonymous):

thank you all very much! I understand the process much better now.